# Lemmanaid: Neuro-Symbolic Lemma Conjecturing

### v3

@@ 5 Conclusion
Lemmanaid is a novel neuro-symbolic tool for conjecturing lemmas for mathematical formalization. Lemmanaid outperforms both neural baselines we create and symbolic tool, QuickSpec, on test sets from Isabelle’s HOL library and from its Archive of Formal Proofs. Lemmanaid discovers between 29-39.5% of the gold standard human written lemmas in test sets, 8-15% more lemmas than the neural baseline using the same training setup. Lemmanaid can discover useful lemmas for a wide range of formalization domains in mathematics, computer science, and logic.
We note that our experimental setup most likely under-reports results, as we measure matches with one specific gold-standard lemma. It is entirely possible that Lemmanaid sometimes comes up with a different gold-standard lemma from the same theory, or even additional lemmas that are valid and useful but not present in the existing formalization. Our evaluation is a first step towards demonstrating the usefulness of neuro-symbolic conjecturing for proof assistants. So far we have not yet explored its full potential. For instance, for future work Lemmanaid could easily be placed in a workflow utilizing the many tools available in Isabelle/HOL for e.g. counterexample checking and automated proofs: Sledgehammer, simp or even neural proof synthesis methods.
Acknowledgments
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The computation was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Appendix A Appendix
A.1 Template language and symbols
@@
A.2 Lemma Enumeration
Our lemma enumeration strategy is best described through an example. From the Cross_Product_7.thy file from the AFP’s Octonions project, the lemma vector_7 was defined by the user as follows:
lemma vector_7 [simp]:
”(vector [x1,x2,x3,x4,x5,x6,x7] ::(’a::zero)^7)\$1 = x1
”(vector [x1,x2,x3,x4,x5,x6,x7] ::(’a::zero)^7)\$2 = x2
”(vector [x1,x2,x3,x4,x5,x6,x7] ::(’a::zero)^7)\$3 = x3
”(vector [x1,x2,x3,x4,x5,x6,x7] ::(’a::zero)^7)\$4 = x4
”(vector [x1,x2,x3,x4,x5,x6,x7] ::(’a::zero)^7)\$5 = x5
”(vector [x1,x2,x3,x4,x5,x6,x7] ::(’a::zero)^7)\$6 = x6
”(vector [x1,x2,x3,x4,x5,x6,x7] ::(’a::zero)^7)\$7 = x7
Given the way we count lemmas (or what we consider a datapoint in the entirety of the project), we consider this to be 7 lemmas. This stems from wanting our definition of a lemma to have one lemma object and one template associated with it. Isabelle differentiates between a “singleton” fact and an “indexed” fact, so we treat each as a singleton fact. In this example, these 7 lemmas are clearly very similar, but there is nothing enforcing this similarity, except perhaps conventions of Isabelle users. Furthermore, Isabelle allows one to also define multiple lemmas at the same time using a “lemmas” keyword where conventionally the lemmas are still similar but less similar than indexed facts typically are. Because this problem (one command, multiple facts) occurs a number of times across theories, we decided to treat every fact separately.
A.3 Projects in AFP-test
For the formalizations included in our test set AFP-test, Table 5 shows the the formalization topic chosen by the project authors, the number of gold-standard lemmas from each project and the lemma success rates (in percentages) of Lemmanaid, neural-only lemma prediction, and their combination. The results shown are the ensembles of the different variants of Lemmanaid and the neural baselines trained on HOL+AFP-train.
AFP entry Topic(s) Lemmas Lemmanaid Neural Combined
ConcurrentHOL CS/Concurrency 481 34.10% 24.53% 37.63%
Sumcheck CS/Security 62 56.45% 43.55% 58.06%
Protocol CS/Algorithms
Broadcast_Psi CS/Concurrency 145 64.14% 40.69% 66.21%
AutoCorres2 CS/PL 11638 33.17% 16.66% 34.96%
CS/Semantics & reasoning
Tools
Substitutions Logic/Rewriting 55 54.55% 25.45% 56.36%
Lambda-Free
Doob_Convergence Math/Prob. theory 2 0.00 0.00 0.00
Orient_Rewrite Logic/Rewriting 148 47.97% 38.51% 50.00%
Rule_Undecidable
Uncertainty Math/Physics 3 0.00 0.00 0.00
Principle
Derandomization CS/Algorithms 15 46.67% 33.33% 46.67%
with Conditional
Expectations
Verified QBF CS/Algorithms 171 44.44% 39.18% 49.71%
Solving Logic/General logic
IMP CS/PL 50 34.00% 34.00% 36%.00
Noninterference CS/Security
Actuarial Math/Games & econ. 168 0.00 0.00 0.00
Mathematics
LL(1) Parser CS/Algorithms 205 36.59% 27.32% 37.56%
Generator CS/PL
Schönhage-Strassen CS/ALgorithms 128 33.59% 27.34% 37.50%
Multiplication Math/Algebra
Isabelle_DOF CS/Semantics & reasoning 41 48.78% 31.71% 53.66%
Interval Analysis Math/Analysis 694 50.29% 36.89% 52.74%
MFOTL Checker CS/Data mgmt systems 151 43.05% 30.46% 47.02%
CS/Algorithms
Logic/General logic
Decomposition of Math/Algebra 1 0.00 0.00 0.00
totally ordered hoops
Approximate Model CS/Algorithms 112 16.96% 18.75% 21.43%
Counting
Wieferich-Kempner Math/Number theory 8 0.00 0.00 0.00
Theorem
PNT_with Math/Number theory 155 28.39% 21.29% 32.26%
Remainder
Continued Fractions Math/Analysis 428 32.71% 29.67% 37.15%
CondNormReasHOL Logic/Phil. aspects 34 11.76% 2.94% 11.76%
Logic/General logic
Region Quadtrees CS/Data structures 182 35.71% 34.07% 41.21%
Karatsuba CS/Algorithms 431 48.03% 30.63% 51.97%
Pick’s Theorem Math/Geometry 334 14.67% 14.07% 16.77%
Kummer Congruence Math/Number theory 120 34.17% 20.00% 37.50%
A.4 Computing Resources and Details for Replication
We run the majority of our training and evaluation on 8 NVIDIA RTX 2080 Ti (11GB VRAM). We use 8-bit quantization when loading models for both fine-tuning and inference. We train models with DistributedDataParallel (DDP) for 12 epochs and $48e-$ 4 learning rate. We use a maximum sequence length of 1024 with 200 tokens reserved for the output. We use an effective batch size 16.
A.5 Results found using greedy decoding vs beam search
@@
Lemmanaid (types + defs) 27.9% 15.0% 40.9% 25.0% 19.5% 25.7%
Lemmanaid (types) 26.0% 13.6% 44.9% 26.5% 18.4% 28.0%
Lemmanaid (defs) 24.2% 7.3% 36.0% 22.0% 8.9% 26.3%
Neural (types + defs) 24.0% 9.7% 30.3% 20.8% 10.1% 30.6%
Neural (types) 23.4% 10.4% 35.7% 21.7% 11.0% 36.0%
Neural (defs) 21.3% 6.4% 26.0% 20.6% 6.8% 29.4%
Lemmanaid Combined 36.2% 20.5% 55.4% 35.4% 24.7% 43.4%
Neural Combined 31.9% 15.1% 42.3% 30.3% 15.5% 40.3%
Combined 41.1% 25.0% 58.9% 41.0% 29.0% 60.3%
@@
Lemmanaid (types + defs) 37.1% 21.6% 50.0% 36.2% 27.3% 44.9%
Lemmanaid (types) 33.4% 22.5% 56.6% 36.4% 24.6% 49.7%
Lemmanaid (defs) 31.3% 10.4% 38.6% 31.2% 13.4% 36.0%
Neural (types + defs) 25.7% 10.4% 23.7% 25.7 % 13.4% 31.4%
Neural (types) 23.6% 13.8% 40.0% 22.9% 11.7% 41.1%
Neural (defs) 21.5% 5.3% 21.1% 25.0% 9.3% 32.0%
Lemmanaid Combined 45.4% 28.5% 67.1% 46.4% 33.2% 64.6%
Neural Combined 37.8% 17.8% 47.1% 36.6% 19.3% 55.4%
Combined 49.3% 30.9% 70.9% 50.4% 36.1% 72.9%
@@
We break down the results from Table 1 into Table 6 and Table 7 to show the performance of different decoding strategies. The results in Table 6 are for greedy decoding, which selects the token with the highest probability at each generation step. The results in Table 7 are for beam search, which obtains multiple high-probability predictions for a given model. We see that beam search helps to improve results for both Lemmanaid and the neural baseline ensembles. However, Lemmanaid is still performant in a greedy setting, showing that the approach can achieve good results with less inference budget.
A.6 Partial QuickSpec output on Octonions

### v5

@@ 5 Conclusion
Lemmanaid is a novel neuro-symbolic bottom-up conjecturing tool for the Isabelle proof assistant. Lemmanaid works by (neurally) learning analogical families of conjectures, represented as templates which are symbolically instantiated in new theories. Lemmanaid outperforms both neural baselines and a symbolic tool, QuickSpec, on test sets from Isabelle’s HOL library and from its Archive of Formal Proofs, discovering 50% of gold standard lemmas for HOL-test and 28% for AFP-test, using deepseek-6.7b as backend. Prompting strategies including both type information and definitions seem to work slightly better for the deepseek backends, while only including type information was favorable for Llama-3.2, although differences are small. Ensembling the different strategies further improves results, showing that the different approaches, including the neural baseline, discovers slightly different conjectures.
This work is a first step toward demonstrating the usability of neuro-symbolic analogical conjecturing for proof assistants on a realistic selection of theories from real formalizations in computer science, mathematics and logic, going beyond benchmark theories from high-school and undergraduate mathematics competitions. In the future, Lemmanaid could be integrated in a workflow supporting for example auto-formalization, especially complementing works on auto-formalization of mathematical definitions.
Limitations
Due to limitations in computational resources, we have evaluated the Lemmanaid framework with smaller LLM backends, ranging between 1B to 6.7B, for two different model families, DeepSeek and Llama 3. Each instance of Lemmanaid requires fine-tuning both the template-generation model, and a direct neural conjecturing model. This in turn, was done for two (or three) different prompting strategies (see e.g. table 1), including either both type and function definitions, or only one of them. Further prompting experiments, and wider beam searches were beyond the scope of our computational budget. Furthermore, for the largest model (Deepseek-coder-6.7b) we suspect additional computational resources might have revealed slightly better hyper-parameter settings, which with additional fine-tuning could have improved those results further.
We also conducted some brief preliminary experiments to compare with commercial LLMs, which however was difficult to do in a controlled and fair setting to compare. We found that the commercial models tended to either expose suspected test-data leakage, re-producing the target Isabelle-formalization almost verbatim, or didn’t succeed at all, in line with results in Johansson and Smallbone (2023).
Finally, as highlighted in §4.3, Lemmanaid’s symbolic engine could be further improved to prune the number of symmetric instantiations, leading to duplicates and equivalent conjectures. This happens especially in the presence of templates with several logic combinators (or, and). In addition, also described in in §4.3, the conjectures generated by Lemmanaid has only been attempted to be proved by the fully automated tactic Sledgehammer. While powerful, any conjecture requiring a more specialized tactic-based proof remain open. In this paper, our focus was a comparison across a wide range of formalizations against their gold-standard lemmas, why we leave deeper analysis and interactive proof attempts of open conjectures in each theory, as future work.
Appendix A Template language and symbols
@@
Appendix B Computing Resources and Details for Replication
We run the majority of our training and evaluation on 8 NVIDIA RTX 2080 Ti (11GB VRAM). The Deepseek 6.7B models are trained and evaluated on 4 NVIDIA Tesla A40s, each with 48GB of VRAM. We use 8-bit quantization when loading models for both fine-tuning and inference. We train models with DistributedDataParallel (DDP) for 12 epochs and 8e48e-4 learning rate. We use a maximum sequence length of 1024 with 200 tokens reserved for the output. We use an effective batch size 16.
Appendix C Template instantiation by an LLM vs Lemmanaid’s symbolic engine
As discussed in Section 3.1, Lemmanaid instantiates its predicted templates symbolically. As an alternative approach, we also experimented with trining an LLM (DeepSeek-coder-1.3b) to instantiate templates. We define the instantiation rate as measuring the percentage of instantiation tasks that the respective methods can perform successfully.
We observe that Lemmanaid’s symbolic engine has a much greater instantiation rate (89.1%) than that of the LLM (66.9%), see Table 3. However, interestingly, the LLM recovers some lemmas that we symbolically fail to recover, so when we consider lemmas recovered either symbolically or neurally, the rate is 92.2%.
Method Instantiation Rate
LLM 66.9%
Lemmanaid 89.1%
Combined 92.2%
Appendix D Results found using greedy decoding vs beam search
@@
Lemmanaid (types + defs) 28.6% 14.5% 47.1% 29.0% 19.6% 51.4%
Lemmanaid (types) 27.2% 14.8% 48.0% 24.3% 19.6% 42.6%
Lemmanaid (defs) 23.6% 4.7% 44.3% 24.2% 8.0% 53.1%
Neural (types + defs) 22.8% 7.0% 29.1% 25.5% 11.0% 37.1%
Neural (types) 19.9% 7.7% 26.6% 23.2% 9.7% 41.7%
Neural (defs) 22.0% 3.4% 31.4% 20.3% 4.7% 36.3%
Lemmanaid Combined 37.1% 20.6% 58.0% 37.3% 26.8% 61.4%
Neural Combined 31.7% 11.3% 44.6% 32.8% 15.3% 50.6%
Combined 42.5% 22.7% 62.0% 42.9% 29.6% 66.6%
Deepseek-coder-6.7b
Lemmanaid (types+defs) 34.7% 19.6% 60.6% 36.3% 9.8% 62.0%
Lemmanaid (types) 31.5% 17.9% 54.0% 34.3% 18.8% 45.7%
Neural (types+defs) 29.7% 11.1% 46.9% 15.8% 3.1% 33.4%
Neural (types) 28.1% 10.7% 42.9% 8.6% 14.1% 46.0%
Lemmanaid Combined 40.1% 22.9% 66.0% 42.2% 21.3% 69.1%
Neural Combined 34.7% 14.9% 50.9% 17.6% 15.2% 52.3%
Combined 45.0% 25.7% 70.3% 44.1% 26.2% 74.3%
Llama-3.2-1b
Lemmanaid (types+defs) 16.2% 9.0% 41.4% 16.8% 9.4% 20.0%
Lemmanaid (types) 17.8% 9.1% 43.4% 19.3% 11.1% 22.9%
Neural (types+defs) 11.5% 3.4% 15.4% 9.5% 4.1% 36.3%
Neural (types) 11.2% 3.8% 21.4% 10.1% 4.3% 34.6%
Lemmanaid Combined 23.2% 13.5% 49.7% 24.2% 18.8% 52.6%
Neural Combined 16.3% 5.1% 30.0% 14.5% 7.3% 18.3%
Combined 27.4% 14.7% 51.7% 27.5% 20.2% 54.0%
@@
Lemmanaid (types + defs) 37.4% 18.7% 52.6% 39.8% 26.2% 63.7%
Lemmanaid (types) 35.6% 17.8% 49.7% 33.3% 25.7% 52.9%
Lemmanaid (defs) 31.6% 6.6% 48.6% 33.4% 12.1% 58.3%
Neural (types + defs) 28.2% 9.1% 31.4% 27.0 % 12.3% 43.4%
Neural (types) 31.9% 10.0% 42.3% 27.0% 11.7% 44.0%
Neural (defs) 29.2% 5.3% 41.4% 23.3% 6.4% 40.6%
Lemmanaid Combined 46.4% 25.0% 66.0% 48.0% 33.8% 74.6%
Neural Combined 40.6% 14.4% 52.9% 36.1% 17.3% 56.3%
Combined 52.0% 27.0% 69.4% 51.4% 35.8% 78.9%
Deepseek-coder-6.7b
Lemmanaid (types+defs) 46.0% 25.8% 68.0% 49.9% 27.4% 76.6%
Lemmanaid (types) 42.9% 16.5% 66.6% 47.2% 28.2% 79.1%
Neural (types+defs) 40.9% 16.8% 52.6% 21.8% 20.0% 61.7%
Neural (types) 37.1% 14.7% 51.1% 36.7% 18.9% 52.9%
Lemmanaid Combined 51.6% 28.4% 76.0% 53.9% 34.1% 83.1%
Neural Combined 45.3% 20.0% 58.9% 41.7% 25.1% 65.1%
Combined 56.4% 30.9% 79.4% 57.4% 38.9% 84.6%
Llama-3.2-1b
Lemmanaid (types+defs) 31.1% 17.8% 57.4% 27.2% 21.2% 46.6%
Lemmanaid (types) 33.5% 18.4% 59.7% 34.0% 23.1% 60.6%
Neural (types+defs) 21.6% 7.0% 32.0% 18.0% 9.7% 54.3%
Neural (types) 21.7% 7.8% 41.1% 14.3% 9.3% 56.9%
Lemmanaid Combined 38.6% 20.2% 62.6% 37.4% 29.0% 64.3%
Neural Combined 27.3% 8.2% 46.0% 21.9% 11.5% 36.3%
Combined 41.7% 21.2% 63.4% 39.7% 30.5% 66.0%
@@
We break down the results from Table 1 into Table 4 and Table 5 to show the performance of different decoding strategies. The results in Table 4 are for greedy decoding, which selects the token with the highest probability at each generation step. The results in Table 5 are for beam search, which obtains multiple high-probability predictions for a given model. We see that beam search helps to improve results for both Lemmanaid and the neural baseline ensembles. However, Lemmanaid is still performant in a greedy setting, showing that the approach can achieve good results with a smaller inference budget.
Appendix E Breakdown of Results per Theory
E.1 HOL-test
For the lemmas included in our test set HOL-test, Table 6 shows the number of test lemmas from each session and the lemma success rates (in percentages) of Lemmanaid, neural-only lemma prediction, and their combination. The results shown are the ensembles of the different variants of Lemmanaid and the neural baselines, trained on HOL+AFP-train using DeepSeek-coder-6.7b. We created a file-wise split of the HOL library.
For the 29 different sessions, we see that the Lemmanaid ensemble outperforms the neural ensemble on 25 of them, while the neural ensemble performs better in one case, and in the remaining three both methods get the same rate of accurate predictions (0% in two of the three). On average the success rate of the Lemmanaid ensemble is 10.34% higher than that of the neural ensemble. Their combined success rate is on average 3.00% higher than that of the Lemmanaid ensemble, showing that the approaches are somewhat complementary but mostly overlapping.
A clear outlier is the Eisbach session where both methods are completely unsuccessful, this is because the lemmas in that set are all simple example lemmas used to demonstrate the use of the Eisbach proof method language Matichuk et al. (2016). These examples  [^8] https://isabelle.in.tum.de/library/HOL/HOL-Eisbach/Examples.html only contain basic logic symbols which we discard from our input symbol set, so our models don’t have any input to base their predictions on.
Table 6: Results for the different sessions of the HOL library.
HOL session Lemmas Lemmanaid Neural Combined
Algebra 49 73.47% 67.35% 81.63%
Analysis 1009 41.72% 28.84% 44.80%
Auth 350 61.71% 60.00% 64.29%
Bali 346 40.46% 21.10% 41.62%
Cardinals 93 46.24% 34.41% 46.24%
Computational_Algebra 50 66.00% 62.00% 66.00%
Data_Structures 91 74.73% 70.33% 76.92%
Datatype_Examples 145 53.10% 23.45% 55.17%
Decision_Procs 240 58.75% 47.08% 60.42%
Eisbach 23 0.00% 0.00% 0.00%
ex 75 41.33% 37.33% 42.67%
Hoare 15 80.00% 66.67% 80.00%
HOLCF 93 61.29% 47.31% 63.44%
IMP 284 89.08% 88.73% 95.07%
Imperative_HOL 61 40.98% 27.87% 45.90%
IMPP 26 50.00% 26.92% 53.85%
IOA 20 45.00% 35.00% 50.00%
Lattice 55 61.82% 54.55% 69.09%
Library 1054 54.74% 39.47% 58.25%
Matrix_LP 25 24.00% 20.00% 24.00%
Metis_Examples 1 0.00% 0.00% 0.00%
MicroJava 178 47.75% 34.27% 51.12%
Nominal 302 81.79% 61.26% 84.44%
Nonstandard_Analysis 36 75.00% 72.22% 80.56%
Number_Theory 47 61.70% 61.70% 72.52%
Quotient_Examples 18 55.56% 38.89% 55.56%
Proofs 6 33.33% 16.67% 33.33%
TLA 7 57.14% 28.57% 57.14%
UNITY 41 14.63% 19.51% 24.39%
E.2 AFP-test
For the formalizations included in our test set AFP-test, Table 7 shows the formalization topic chosen by the project authors, the number of gold-standard lemmas from each project and the lemma success rates (in percentages) of Lemmanaid, neural-only lemma prediction, and their combination. The results shown are the ensembles of the different variants of Lemmanaid and the neural baselines, trained on HOL+AFP-train using DeepSeek-coder-6.7b.
We can see that out of the 27 tested formalization projects, the Lemmanaid ensemble outperforms the neural ensemble on 19 of them, while the neural ensemble performs better for two, and in the remaining 6 cases the performance is equally good (0% in five out of six cases). Note that most of the projects with a success rate of 0% contain a small number of target lemmas (less than 10), with the outlier being the Actuarial Mathematics formalization. In this case, the syntax used in the formalization meant that none of the predicted lemmas were considered syntactically equivalent, however a more thorough evaluation would have shown some of them to be semantically equivalent to the target lemmas. On average the Lemmanaid ensemble has a 7.18% better performance than the neural one. Their combined success rate is on average 3.52% higher than the one for the Lemmanaid ensemble.
AFP entry Topic(s) Lemmas Lemmanaid Neural Combined
ConcurrentHOL CS/Concurrency 481 39.71% 28.27% 43.24%
Sumcheck CS/Security 62 67.74% 62.90% 74.19%
Protocol CS/Algorithms
Broadcast_Psi CS/Concurrency 145 82.07% 69.66% 85.52%
AutoCorres2 CS/PL 12721 34.72% 22.48% 37.86%
CS/Semantics & reasoning
Tools
Substitutions Logic/Rewriting 55 50.91% 21.82% 52.73%
Lambda-Free
Doob_Convergence Math/Prob. theory 2 0.00% 0.00% 0.00%
Orient_Rewrite Logic/Rewriting 148 60.14% 42.57% 62.16%
Rule_Undecidable
Uncertainty Math/Physics 3 0.00% 0.00% 0.00%
Principle
Derandomization CS/Algorithms 15 40.00% 40.00% 40.00%
with Conditional
Expectations
Verified QBF CS/Algorithms 171 53.80% 43.86% 57.31%
Solving Logic/General logic
IMP CS/PL 50 40.00% 38.00% 42.00%
Noninterference CS/Security
Actuarial Math/Games & econ. 168 0.00% 0.00% 0.00%
Mathematics
LL(1) Parser CS/Algorithms 205 38.05% 46.83% 55.61%
Generator CS/PL
Schönhage-Strassen CS/ALgorithms 128 41.41% 26.56% 43.75%
Multiplication Math/Algebra
Isabelle_DOF CS/Semantics & reasoning 41 53.66% 34.15% 53.66%
Interval Analysis Math/Analysis 694 55.19% 46.83% 58.50%
MFOTL Checker CS/Data mgmt systems 2081 14.22% 24.89% 30.71 %
CS/Algorithms
Logic/General logic
Decomposition of Math/Algebra 1 0.00% 0.00% 0.00%
totally ordered hoops
Approximate Model CS/Algorithms 112 32.14% 17.86% 34.82%
Counting
Wieferich-Kempner Math/Number theory 8 0.00% 0.00% 0.00%
Theorem
PNT_with Math/Number theory 155 35.48% 25.81% 40.00%
Remainder
Continued Fractions Math/Analysis 428 37.62% 35.75% 43.69%
CondNormReasHOL Logic/Phil. aspects 34 11.76% 5.88% 11.76%
Logic/General logic
Region Quadtrees CS/Data structures 182 41.76% 34.07% 46.15%
Karatsuba CS/Algorithms 431 50.58% 32.25% 55.22%
Pick’s Theorem Math/Geometry 334 19.46% 14.37% 22.46%
Kummer Congruence Math/Number theory 120 34.17% 25.83% 38.33%
Appendix F Examples of Generated Conjectures
The analysis presented in section 4.3 divides the set of conjectures generated by Lemmanaid into true, false, and open conjectures. Here, we present some examples of conjectures in each category.
True Statements
If Sledgehammer succeeds in finding a proof of a given conjecture, it’s considered a true statement. True statements included both gold-standard lemmas formalized by experts and other novel true statements not previously formalized. The following two conjectures were discovered by Lemmanaid: the first lemma is identical to the gold-standard lemma totient_imp_primetotient\_imp\_prime and the second one doesn’t match any formalized lemmas.
lemma.
$totient\ x_{1}=x_{1}-1\Longrightarrow 0<x_{1}\Longrightarrow prime\ x_{1}$"
lemma.
$1\leq x_{1}\Longrightarrow 1\leq totient\ x_{1}$"
False Statements
Conjectures for which Nitpick or Quickcheck can find a counter-example are considered false statements. For the conjecture below, Quickcheck finds a counter-example when $x_{1}=2$.
lemma.
$2\leq x_{1}\Longrightarrow even\ (totient\ x_{1})$"
Open Conjectures
Lemmanaid generates a number of conjectures which can’t be automatically categorized. This is expected due to the incompleteness of the automated reasoning tools we use. These conjectures require further complex reasoning to decide if that are true or false. One example of an open conjecture generated by Lemmanaid is the monotonicity of Euler’s function. Nitpick and Quickcheck fail to find a counterexample such as $x_{1}=21,x_{2}=22$.
lemma.
$x_{1}\leq x_{2}\Longrightarrow totient\ x_{1}\leq totient\ x_{2}$"
Appendix G Additional Test Set Statistics
Statistics on HOL-test, AFP-test, and Octonions, and the successes of Lemmanaid and neural baselines on each dataset are shown in Table 8. We count the number of lemmas in each test set (Lemmas), how many of them are equational (Eq), and look into the length (in characters) of lemmas and their templates as well as the number of different symbols that appear. Lemmanaid and neural models here are trained on HOL-train and use greedy decoding, using DeepSeek-coder-6.7b for both the neural prediction and the neural engine of Lemmanaid.
Table 8: Number of lemmas in each test set, and how many were discovered by Lemmanaid and the corresponding neural method (using DeepSeek-6.7b and greedy decoding), including how many were equational. The table also shows information about template and lemma lengths (number of characters) and the number of different symbols appearing.
Lemmas Eq Template Length Lemma Length # Symbols
Min 25% Med 75% Max Min 25% Med 75% Max Mean Max
HOL-test 4740 2242 9 49 82 127 9014 7 57 91 150 10454 4.2 100
Lemmanaid 1643 885 11 37 55 85 300 10 45 69 100 368 3.3 15
Neural 1407 750 11 37 55 85 433 11 45 68 98 390 3.3 15
AFP-test 18975 13202 7 39 63 130 10492 7 58 80 141 7530 4.1 31
Lemmanaid 3712 2989 7 25 31 45 220 12 39 55 73 286 2.8 14
Neural 2111 1526 7 23 33 45 184 10 36 52 71 253 2.8 12
Octonions 350 262 17 29 49 75 688 11 33 55 79 391 3.8 21
Lemmanaid 212 152 17 23 43 67 121 11 29 47 65 133 2.7 11
Neural 164 105 19 23 49 67 269 15 25 47 66 279 2.7 11.0
Appendix H Partial QuickSpec output on Octonions